Hi! I need help solving
(x^-2y^-2/xz)/(xy/y^-2z^3)
thanks in advance for your help!

Answer by tanimachatterjee(60)

(x^-2y^-2/xz)/(xy/y^-2z^3)
= x^(-2-1)y^(-2)z^(-1)/xy^(1+2)z^(-3)
=x^(-3)y^(-2)z^(-1)/xy^(3)z^(-3)
=x^(-3-1)y^(-2-3)z^(-1+3)
=x^(-4)y^(-5)z^2